Hello i need help with for proportion problems if it's possible,the first question is solve the proportion,4over3.2=2overx.The second one is solve the proportion,zover3=z+9over12.The next one is solve the proportion 16over4d+2=8over3.And the last one is find the probability that the name of a month of the year,chosen at random,begins with a consonant and that question the last one i really don't understand at all thanks for the help
Answer by checkley75(3666)

4/3.2=2/x cross multiply
4x=6.4
x=6.4/4
x=1.6 answer
----------------------------
z/3=(z+9)/12
12z=3z+27
12z-3z=27
9z=27
z=27/9
z=3 answer
-------------------------------
16/(4d+2)=8/3
8(4d+2)=16*3
32d+16=48
32d=48-16
32d=32
d=32/32
d=1 answer